# Satisfaction-Aware Incentive Scheme for Federated Learning in Industrial Metaverse: DRL-Based Stackbelberg Game Approach

### v1

@@ V-B Stackelberg Game Analysis
VI DRL-based Stackelberg Game Approach
@@ VI-B MADDPG Algorithm
The process of the DRL-based Stackelberg game is illustrated in Fig. 2, where the server serves as the leader and nodes act as the followers. In each training cycle, the server agent observes the state $o^{t}_{server}$ and determines the action $r^{t}_{server}$, while the node agent observes the state $o^{t}_{node}$ and determines the action $\theta^{t}_{node}$. Subsequently, the current state transitions to the next, and the agents receives the reward. The detailed components of the DRL controller for each agent are shown in Fig. 3. The replay buffer is used to store transitions, including the current state, action, reward, and next state, collected during interactions with the environment. These stored transitions are sampled in batches to decorrelate sequential data and stabilize the training process. The actor and critic network comprising three fully connected layers. The actor network takes the current state as input and outputs the corresponding action by generating a policy. The critic network evaluates the action taken by the actor network and provides a value estimation to guide policy improvement. Both the actor and critic networks are updated using two separate optimization modules. The policy optimizer updates the parameters of the actor network based on the policy gradient, while the value optimizer minimizes the temporal difference error to refine the critic network’s value estimations.
Algorithm 1 MADDPG-based Solution for Stackelberg Game
@@
VII Simulation Results
To verify the impact of the proposed satisfaction-aware incentive scheme on FL in the industrial Metaverse, this paper conducts experiments using Python 3.7 and TensorFlow 1.15, and evaluates the performance of the scheme. The value ranges of the simulation parameters are shown in Table II. We take image classification tasks as an example of industrial Metaverse applications. In the physical space, the robots perform image classification tasks without loss of generality. We use the MNIST dataset, consisting of 60,000 samples for the training set and 10,000 samples for the test set. To evaluate the performance of the proposed scheme, we compare it with four reinforcement learning algorithms: MADDPG [33], MAPPO [40], MASAC [41], and MADQN [42], which are DDPG algorithm, PPO algorithm, SAC algorithm, and DQN algorithm in multi-agent environment.
@@ VII-C Solution Comparison
Figure 9: FL accuracy for different number of selected nodes.
@@ VIII Conclusion
In this paper, we have designed a satisfaction-aware FL incentive scheme for the industrial Metaverse. The scheme integrates a satisfaction function that incorporates data size, AoI, and latency into the utility function and formulates this utility optimization problem as a two-stage Stackelberg game. We have employed a DRL approach to learn the equilibrium of the Stackelberg game, with the goal of maximizing the server’s utility by identifying the optimal equilibrium. Experimental results have demonstrated that, compared to traditional methods, the scheme effectively balances model quality and update latency through efficient resource allocation, enhances FL performance, ensures real-time performance and high efficiency of FL in the Metaverse, and better addresses the needs of the industrial Metaverse. In the future, we will extend our meta-computing framework to incorporate efficient asynchronous FL, with the aim of enhancing learning efficiency in the industrial Metaverse.

### v2

@@ V-B Stackelberg Game Analysis
VI DRL-based Stackelberg Game
@@ VI-B MADDPG Algorithm
The process of the MADDPG-based Stackelberg game is illustrated in Fig. 2, where the server serves as the leader and nodes act as the followers. In each training cycle, the server agent observes the state $o^{t}_{server}$ and determines the action $r^{t}_{server}$, while the node agent observes the state $o^{t}_{node}$ and determines the action $\theta^{t}_{node}$. Subsequently, the current state transitions to the next, and the agents receives the reward. The detailed components of the DRL controller for each agent are shown in Fig. 3. The replay buffer is used to store transitions, including the current state, action, reward, and next state, collected during interactions with the environment. These stored transitions are sampled in batches to decorrelate sequential data and stabilize the training process. The actor and critic network comprising three fully connected layers. The actor network takes the current state as input and outputs the corresponding action by generating a policy. The critic network evaluates the action taken by the actor network and provides a value estimation to guide policy improvement. Both the actor and critic networks are updated using two separate optimization modules. The policy optimizer updates the parameters of the actor network based on the policy gradient, while the value optimizer minimizes the temporal difference error to refine the critic network’s value estimations.
Algorithm 1 MADDPG-based Stackelberg Game
@@
VII Simulation Results
To verify the impact of the proposed satisfaction-aware incentive scheme on FL in the IIoT, this paper conducts experiments using Python 3.7 and TensorFlow 1.15, and evaluates the performance of the scheme. The value ranges of the simulation parameters are shown in Table II. We take image classification tasks as an example of IIoT applications. In the physical space, the robots perform image classification tasks without loss of generality. We use the MNIST dataset, consisting of 60,000 samples for the training set and 10,000 samples for the test set. To evaluate the performance of the proposed scheme, we compare it with four reinforcement learning algorithms: MAPPO [35], MASAC [36], and MADQN [37], which are DDPG algorithm, PPO algorithm, SAC algorithm, and DQN algorithm in multi-agent environment.
@@ VII-B FL Performance Comparison
Figure 9: FL accuracy for different number of selected nodes.
@@ VIII Conclusion
In this paper, we have designed the MEFL framework for IIoT. The framework leverages meta-computing to enhance FL in IIoT. We have developed a satisfaction-aware incentive scheme within MEFL. This scheme integrates a satisfaction function that incorporates data size, AoI, and latency into the utility function, and formulates the utility optimization problem as a two-stage Stackelberg game. We employ a DRL-based approach to learn the equilibrium of the Stackelberg game, aiming to maximize the server’s utility by identifying the optimal equilibrium. Experimental results demonstrate that, compared to traditional methods, the scheme effectively balances model quality and update latency through efficient resource allocation, enhances FL performance, ensures real-time capability and high efficiency of FL in IIoT, and better addresses the specific needs of industrial scenarios. In future work, we will extend our meta-computing framework to incorporate efficient asynchronous FL, with the aim of further improving learning efficiency in IIoT.